Circular No. 6684

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1997 L2 (SOHO) C. St. Cyr, Naval Research Laboratory, reports, for the SOHO-LASCO

Consortium (see IAUC 6669), the discovery of a faint comet in SOHO-LASCO data by Shane Stezelberger last week. This comet was moving away from the sun, and there is again some inconsistency between the data obtained with the C2 and the C3 coronagraphs. The comet has no obvious tail. Positions (apparent solar distance and position angle) have been measured by D. Lewis (University of Birmingham), S. Paswaters, St. Cyr and Stezelberger. These positions have been reduced to the J2000.0 earth-equatorial system by G. V. Williams and are given in detail on MPEC 1997-M06. The following is the approximate position on the first C2 image:

1997 UT  α_{2000}  δ_{2000}  m_{1}
June 10.592  5^{h}15.^{m}1  +23^{o}48′  5

This comet is clearly not a Kreutz sungrazer, and the rough parabolic orbital elements below have been determined by the undersigned. The comet is likely to be considerably fainter than suggested in the ephemeris.

T = 1997 June 10.98 TT  ω = 66.^{o}39
Ω = 72.79 }2000.0
q = 0.0430 AU  i = 70.79
1997TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
June 20 6^{h}07.^{m}53 + 5^{o}23.^{′}8 0.665 0.438 18.^{o}3 133.^{o}2 13.5
21 6 13.15 + 2 34.0 0.650 0.472 21.2 129.0 13.8
22 6 18.91 − 0 18.0 0.639 0.505 24.1 124.9 14.1
23 6 24.83 − 3 11.3 0.629 0.538 27.0 121.0 14.3
24 6 30.90 − 6 05.1 0.622 0.569 29.9 117.1 14.5
25 6 37.12 − 8 58.3 0.617 0.600 32.8 113.3 14.7
26 6 43.48 −11 50.0 0.614 0.630 35.7 109.7 14.9
27 6 49.99 −14 39.3 0.613 0.659 38.5 106.1 15.1
28 6 56.63 −17 25.3 0.614 0.688 41.3 102.6 15.3
29 7 03.40 −20 07.1 0.617 0.716 44.0  99.2 15.5
30 7 10.29 −22 43.9 0.621 0.743 46.7  95.9 15.7
July 1 7 17.31 −25 15.3 0.627 0.771 49.2  92.8 15.9

CORRIGENDUM SN 1997by (cf. IAUC 6639, 6640, 6643) is in IC 4252. On IAUC 6640,

lines 7–8, for parent galaxy Abell 1736 read parent galaxy in Abell 1736

1997 June 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden